OpenStudy (anonymous):

Which expression is the simplified form of : the square root of 338w^18??

OpenStudy (anonymous):

\[\sqrt{338w ^{18}}=\sqrt{2*169*w ^{2*9}}=13w ^{9} \sqrt{2}\]